please reduce the fraction. (-20)(t^5)(u^2)(v^3) / (48)(t^7)(u^4)(v)
thank you!!

Answer by smik(40)

Good question!


(-20)(t^5)(u^2)(v^3) / (48)(t^7)(u^4)(v)


considering its the division of these variables, we subtract the power of the variable in the numerator, by the power of the same variable in the denominator. E.g. So: t^5 / t^7 = t^(-2). Furthermore both -20 and 48 are divisible by 4, giving us -5 and 12 respectively.


Therefore the above comes down to:


(-5)(t^-2)(u^-2)(v^2) / (12)


Now the final step is to shift all the negative exponents in the numerator down to the denominator. E.g. 9^-2 = 1/(9^2).


So we get (as a final answer):


(-5)(v^2) / (12)(t^2)(u^2)